Ella Ltd provides its employees with long service leave (LSL) entitlements of 13 weeks of paid leave for every 10 years of continuous service. As the company has been operating for only 2 years, no employees have become entitled to long service leave. However, the company recognises a provision for long service leave in accordance with AASB 119/IAS 19. The following information is obtained from Ella Ltd’s payroll records and actuarial reports for the employees at 30 June 2020.

Total number of employees

100

Annual salary per employee

$50 000

Number of years (unit credit)

2 years

Number of years until long service leave is expected to be paid

8 years

Probability that long service leave will be taken (proportion of employees expected to stay long enough to become entitled to long service leave)

20%

Yield on 8-year high-quality corporate bonds at 30/6/2020

10%

•        Additional information

–       The estimated annual increase in wages is 2% p.a. for the next 10 years, reflecting expected inflation.

–       The provision for long service leave for the employees at 30 June 2019 was $12  000.

REQUIRED

Provide the numbers for each step and the journal entry in the blanks below  to account for Ella Ltd’s provision for long service leave at 30 June 2020. Round to the nearest $ amount.

--------

Step 1: Estimate the number of employees who are expected to become eligible for long service leave:  

Step 2: Estimate the projected salaries:  

Step 3: Determine the accumulated benefit:  

Step 4: Measure the present value of the accumulated benefit:  

Journal entry:

30/6/2020

Long Service Leave Expense

Dr

   Provision for Long Service Leave

Cr
Answer #1

Step 1: Estimate the number of employees who are expected to become eligible for long service leave:  

No of Employees expected to become eligible = 100*20% = 20 employees

Step 2: Estimate the projected salaries:

Salary = $50000*1.02^7 = $57434 p.a. per employee

Step 3: Determine the accumulated benefit:

Accumulated Benefit = $57434*20 employees*13/52 = $287170

Step 4: Measure the present value of the accumulated benefit:

Present value of the accumulated benefit = $287170/1.1^7 = $147364

Step 4: Current year's provision for long service leave for the employees to be created = $147364 - $12000 = $135364

Journal entry:

30/6/2020

Long Service Leave Expense

Dr

$135364

   Provision for Long Service Leave

Cr

$135364